OpenStudy (anonymous):

35 feet is what percent of 244 feet?

OpenStudy (mathlegend):

Did you ever learn... this.. \[\frac{ is }{ of } = \frac{ percent }{ 100 }\]

OpenStudy (mathlegend):

35 feet is of 244 feet See the two key words there that you can now fill in? @Pash007

OpenStudy (mathlegend):

Can you fill in the proportion @Pash007 ?

OpenStudy (anonymous):

14.34?

OpenStudy (mathlegend):

Is that the answer you got?

OpenStudy (anonymous):

yes

OpenStudy (mathlegend):

Can you show me your steps please?

OpenStudy (anonymous):

100/244=0.40 x 35= 14.34

OpenStudy (mathlegend):

Okay, great, so does it ask you to round anymore?

OpenStudy (anonymous):

no, but when i placed this in my answer box it says incorrect..

OpenStudy (mathlegend):

Hmm... that is quite interesting...

OpenStudy (mathlegend):

What if you found to the tenths or whole number?

OpenStudy (anonymous):

but, I did find a way to express the answer in the terms its asking for.. 875/61 weird, but Correct...

OpenStudy (mathlegend):

Oh, it wanted a fraction...